# Higher-order triadic percolation on random hypergraphs

### v1

@@ Abstract
In this work we propose a comprehensive theoretical framework combining percolation theory to nonlinear dynamics in order to study hypergraphs with a time-varying giant component. We consider in particular hypergraphs with higher-order triadic interactions that can upregulate or downregulate the hyperedges. Triadic interactions are a general type of signed regulatory interaction that occurs when a third node regulates the interaction between two other nodes. For example, in brain networks, the glia can facilitate or inhibit synaptic interactions between neurons. However, the regulatory interactions may not only occur between regulator nodes and pairwise interactions but also between regulator nodes and higher-order interactions (hyperedges), leading to higher-order triadic interactions. For instance, in biochemical reaction networks, the enzymes regulate the reactions involving multiple reactants. Here we propose and investigate higher-order triadic percolation on hypergraphs showing that the giant component can have a non-trivial dynamics. Specifically, we demonstrate that, under suitable conditions, the order parameter of this percolation problem, i.e., the fraction of nodes in the giant component, undergoes a route to chaos in the universality class of the logistic map. In hierarchical higher-order triadic percolation we extend this paradigm in order to treat hierarchically nested triadic interactions demonstrating the non-trivial effect of their increased combinatorial complexity on the critical phenomena and the dynamical properties of the process. Finally, we consider other generalizations of the model studying the effect of considering interdependencies and node regulation instead of hyperedge regulation. The comprehensive theoretical framework presented here sheds light on possible scenarios for climate networks, biological networks and brain networks, where the hypergraph connectivity changes over time.
@@ I Introduction
Higher-order networks [1, 2, 3, 4, 5, 6] capture the many-body interactions among two or more nodes present in a large variety of complex systems, ranging from the brain to chemical reaction networks. Higher-order networks display a very rich interplay between topology and dynamics [1, 2, 7] which is leading to significant progress in network theory. Notably, higher-order interactions affect synchronization [8, 9, 10, 11, 12], percolation [13, 14, 15, 16, 17, 18, 19, 20], epidemic spreading [21, 22, 23], and random walks and diffusion [24, 25].
Triadic interactions are a general type of higher-order interactions in which one or more nodes regulate the activity of a link between two other nodes. This type of interaction is widely observed in nature such as brain networks, ecological networks, and biochemical reaction networks [26, 27, 28, 29]. Indeed, in the brain, the glia inhibit or facilitate the synaptic interactions between pairs of neurons; in biochemical reaction networks, the reactions are controlled by increasing or decreasing the activity of enzymes. Recently, there is a growing interest in triadic interactions from theoretical side. Importantly, triadic interactions have been recently shown to give rise to a paradigmatic change in the theory of percolation, according to the recently proposed triadic percolation [13, 30] model. Futhermore, it has been shown that triadic interactions affect learning [31, 32], signal processing and network dynamics [33, 34].
@@
Percolation [36, 37, 38, 39] is one of the most fundamental critical phenomena defined on networks and has been extensively used to characterize the robustness of networks[40]. Percolation on simple networks is characterized by a second order phase transition at which the emergence of the giant component can be observed, when nodes or links are randomly damaged. In the last decades important progress in network theory has demonstrated that more general percolation problems, including k-core percolation [41], and interdependent percolation on multiplex networks [42, 43, 44, 45, 46] can display discontinuous hybrid transitions, higher-order critical points [47] and exotic phase diagrams [48]. In all these problems the size of the giant component is monitored after a (random) damage is inflicted to the nodes or to the links of the network, and after the eventual cascade of failure events triggered by the initial perturbation reaches a steady state.
@@
Triadic percolation [13, 30], however demonstrates that percolation can become a fully-fledged dynamical process in which the giant component never reaches a steady state. If the triadic interactions are defined on top of a random network [13], the giant component becomes in general time-dependent and its size, that defines the order parameter of the model, can undergo a route to chaos in the logistic map universality class [49, 50]. Moreover triadic percolation on spatial networks leads to topological changes in the structure of the giant component generating non-trivial spatio-temporal patterns, as shown in Ref. [30].
@@
In this work we investigate higher-order triadic percolation (HOTP) on hypergraphs. We reveal that on hypergraphs, the HOTP is a fully-fledged dynamical process whose order parameter displays period-doubling and a route to chaos, when both positive and negative regulations are present with non-zero probability. Thus the phase diagram of HOTP reduces in this scenario to an orbit diagram that we hereby prove to be in the universality class of the logistic map.
@@
The critical properties of HOTP are highly non-trivial. Here we characterize them in important limiting cases in which there are exclusively positive or exclusively negative regulatory interactions, characterizing their discontinuous hybrid transitions, period-2 bifurcations, and continuous transitions. When both positive and negative regulations are present, we indicate the conditions for observing unusual orbit diagrams with a reentrant collapsed phase. In this case we observe a phase transition between a collapsed state with a null giant component and an active phase with a non-zero and time-varying giant component while the probability of down-regulating the hyperedges increases.
@@
Triadic interactions can also be hierarchically nested as recognized in the context of ecological networks in Ref. [28]. In order to study the effects of hierarchical regulation, we propose and study hierarchical higher-order triadic percolation (HHOTP). In this scenario, the regulatory interaction between a node and a hyperedge is controlled by other regulatory interactions and so on. We reveal that the dynamics of HHOTP is significantly different from the dynamics of HOTP due to the combinatorial complexity of their HHOTIs. Indeed as long as the hypergraphs contains negative HHOTIs with a non-vanishing probability, HHOTP displays a much richer dynamics than HOTP which can lead to period-doubling and a route to chaos also if no positive HHOTIs are present. Moreover, our theoretical derivation reveals that the route to chaos of HHOTI is no longer in the logistic map universality class.
@@
This work includes an in depth discussion of two other interesting variations of HOTP: the interdependent higher-order triadic percolation (IHOTP), and the higher-order node dynamical percolation model (HONDP). In IHOTP, we adopt the notion of interdependent hypergraph giant component introduced in Ref. [14] which in absence of regulation is already known to display discontinuous hybrid transition. This different nature of the underlying percolation process has dramatic effects on the dynamics of IHOTP. In particular, here we indicate the conditions under which the route-to-chaos in IHOTP is impeded. The study of HONDP allows us to discuss in detail the effect of regulating the nodes instead of the hyperedges revealing important differences with HOTP.
@@
The paper is structured as follows. In Section II, we introduce random hypergraphs with higher-order triadic interactions and with hierarchical triadic interactions; in Sections III and IV we define HOTP and HHOTP and we investigate their critical properties; in Section V, we discuss other generalizations of HOTP, including IHOTP and HONDP; finally, in Section VI we provide the concluding remarks. The paper also includes two Appendices. In the Appendix A we prove that the order parameter of HOTP undergoes a route-to-chaos in the universality class of the logistic map. In Appendix B we discuss HOTP and HONDP in presence of partial regulation, i.e. where hyperedges (for HOTP) or nodes (for HONDP) are regulated with probability $11-\rho_{0}<$ 1.
@@ II.1 Structural hypergraphs
In this work we consider exclusively random hypergraphs with hyperdegree distribution $P(k)$ and hyperedge cardinality distribution $Q(m)$. We denote with $\langle{k}\rangle=\sum_{k}kP(k)$ the average hyperdegree of the nodes of the hypergraph and with $\langle{m}\rangle=\sum_{m}mQ(m)$ the average cardinality of the hyperedges of the hypergraph. We furthermore assume that their factor graph representation is locally tree-like (see for a detailed discussion of this assumption Ref. [14]). This assumption is generally met by the considered random hypergraphs under very general conditions over the distributions $P(k)$ and $Q(m)$, and applies to the case in which $P(k)$ and $Q(m)$ have finite moments as for Poisson distributions. In the following we make use to the hypergraph generating functions $G_{0}(x),G_{1}(x)$ and $G_{1,m}(x)$ defined as
@@
The considered random hypergraphs reduce to networks when all the hyperedges have cardinality $m=2$, (i.e. are edges). In terms of the cardinality distribution $Q(m)$, this happens when $Q(m)=\delta_{m,2}$ where here and in the following $\delta_{x,y}$ indicates the Kronecker delta. In this work we provide general results for arbitrary hypergraphs but if not otherwise explicitly stated, we will perform the numerical analysis and the Monte Carlo simulations on hypergraphs having hyperedges with constant cardinality $m$ and Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$.
@@ II.2.1 Higher-order triadic interactions
Higher-order triadic interactions (HOTIs) occur when one or more nodes regulate a hyperedge. The HOTIs are signed, i.e. they can be associated with a positive or negative regulatory role. For instance, an enzyme can speed up or inhibit a chemical reaction. Here we consider the simple scenario in which the HOTIs occur between the same type of nodes (i.e. we do not distinguish between ”enzyme” and ”metabolite” in the previous example), generalizations in this direction will be considered in future works. In this framework, triadic interactions are encoded by a regulatory bipartite network $\mathcal{G}_{T}(V,E_{H},W)$ formed by the set of nodes $V$ and the set of factor nodes $E_{H}$ where each factor node represents a hyperedge of the structural hypergraph. The signed HOTI are represented by the edges $W$ of the bipartite network $\mathcal{G}_{T}$. Specifically, each edge in $W$ indicates the existence of a (triadic) regulatory interaction from a node in $V$ to a hyperedge in $E_{H}$ (see Fig. 1 (a) and (b)). We call a node a positive (or negative) regulator of a hyperedge if the node regulates the hyperedge positively (or negatively). Note that the sign of regulation is a property of the regulatory interaction but not of a node.
@@ II.2.2 Hierarchical higher-order triadic interactions
Hierarchical higher-order triadic interactions (HHOTI) can be described by a multilayer regulatory network formed by $L$ layers, where each layer $\mu\in\{1,2,\ldots L\}$ is a bipartite network $\mathcal{G}_{T}^{[\mu]}$. The first layer is formed by nodes in $V$ connected to factor nodes in $E_{H}$, i.e. $\mathcal{G}_{T}^{[1]}=(V,E_{H},W^{[1]})$. Thus the first layer reduces to the fundamental HOTI defined in the previous paragraph. The second layer indicates the set of regulatory interactions $W^{[2]}$ of the regulatory interactions in $\mathcal{G}_{T}^{[1]}$, the third layers indicates the regulatory interactions $W^{3}$ of the regulatory interactions in layer 2 and so on (see Fig. 1 (e)). Thus we can describe these further layers with $\mu>1$ as bipartite networks $\mathcal{G}_{T}^{[\mu]}=(V,W^{[\mu-1]},W^{[\mu]})$ fully determining the regulatory interactions $W^{[\mu]}$ between the nodes in $V$ and the factor nodes representing the regulatory interactions $W^{[\mu-1]}$ at the previous layer.
@@
In this work we will focus mostly on numerical results obtained for distributions $\hat{P}_{[\mu]}(\kappa^{\pm})$ independent of the layer $L$, i.e. $\hat{P}_{[\mu]}(\kappa^{\pm})=\hat{P}(\kappa^{\pm})$ with $\hat{P}(\kappa^{\pm})$ being a Poisson distribution with average $c^{\pm}$.
@@ III.1 Higher-order triadic percolation
Here we formulate higher-order triadic percolation (HOTP) that generalizes triadic percolation, proposed for networks in Ref. [13], to hypergraphs. In this framework, nodes are active if they belong to the hypergraph giant component, while hyperedges are upregulated or downregulated according to the activity of their regulator nodes and a stochastic noise. The hypergraph giant component considered here is the largest extensive connected component whose nodes and hyperedges satisfy the following self-consistent and recursive conditions [14].
@@
Given a random hypergraph, at each time $t$, Step 1 implements hyperedge percolation [14] where the probability that a hyperedge is intact is given by $p_{H}^{(t-1)}$. This process determines the fraction of nodes in the hypergraph giant component $R^{(t)}$ at time $t$, given by
@@
where $\hat{S}^{(t)}$ is the probability that starting from a random node and choosing one of its hyperedge at random we reach a hyperedge in the hypergraph giant component. The probability $\hat{S}^{(t)}$ together with the probability $S^{(t)}$ that starting from a random hyperedge and choosing one of its nodes at random we reach a nodes in the hypergraph giant component, obey the following self-consistent set of equations
@@
Thus Eqs. 6 and 7 together with Eq. 5 determine $R^{(t)}$ starting from the knowledge of $p_{H}^{(t)}$ and implements Step 1. Note that this step differs from the Step 1 of standard triadic percolation [13], as here we consider nodes only active if the belong to the hypergraph giant component while in the standard triadic percolation defined on simple networks we define node active if the belong to the standard giant component of their network.
Step 2 requires the formulation of an additional equation determining the probability $p_{H}^{(t)}$ that the hyperedges are intact given the probability $R^{(t)}$ that the nodes are active, (i.e. they belong to the hypergraph giant component at Step 1). This equations reads
@@
where $G_{0}^{\pm}(x)$ are the generating functions defined in Eq. 3.
Note that Eq. 8 is the same equation used in Ref. [13] to define the Step 2 of standard triadic percolation, however here $p_{H}^{(t)}$ indicates the probability that a hyperedge is intact while in standard triadic percolation it indicates the probability that an edge is intact.
From this definition of higher-order triadic percolation it is clear that this dynamical process reduces to standard triadic percolation [13] if $P(m)=\delta_{m,2}$, i.e. if the hypergraph reduces to a network.
Figure 2: Time series (panels (a)-(d)) of the order parameter $R$ of HOTP and their corresponding Cobwebs (panels (e)-(h)) are shown to display a variety of dynamical behavior: a steady state (panels (a) and (e)), period-2 oscillations (panels (b) and (f)), period-4 oscillations (panel (c) and (g)) and chaotic dynamics (panel (d) and (h)). The time series are obtained by Monte Carlo simulations on on hypergraphs of $N=10^{6}$ nodes with HOTIs. The structural hypergraph has hyperedges of fixed cardinality $m=10$ and Poisson hyperdegree distribution $P(k)$ with average $c=10$. The degree distributions $\hat{P}(\kappa^{\pm})$ are Poisson with an average regulatory positive and negative degree $c^{+}=3$ and $c^{-}=5.5$ respectively. In the Cobweb, the blue curves denote function $R=f_{m}(p_{H})$ and the orange curves denote function $p_{H}=g_{p}(R)$ as defined in Eq. 9 and Eq. 10. The timeseries and the Cobweb plots are obtained for $p=0.27$ (panels (a) and (e)), $p=0.35$ (panels (b) and (f)), $p=0.68$ (panels (c) and (g)) and $p=0.88$ (panels (d) and (h)).
Figure 3: Examples of the map $h_{m,p}(R)$ encoding for the dynamics of HOTP when the HOTIs are exclusively positive (panel (a)); when the HOTIs are exclusively negative (panel (b)) and when the hypergraphs include both positive and negative HOTIs with non vanishing probability (panel (c)). The map is unimodal only the scenario (c), thus we can observe a route to chaos of the order parameter only in presence of both positive and negative HOTIs. For all three panels, the structural networks have constant hyperedge cardinality $m=3$ and Poisson hyperdegree distribution with an average $c=4$. Moreover the data ar obtained for $\hat{P}^{\pm}(\kappa)$ with average degree $c^{\pm}$ given by $c^{+}=2,c^{-}=0$ (panel (a)); $c^{+}=0,c^{-}=2$ (panel (b)).; $c^{+}=2,$ $c^{-}=1.5$ (panel (c)).

### v2

@@ Abstract
In this work, we propose a comprehensive theoretical framework combining percolation theory with nonlinear dynamics to study hypergraphs with a time-varying giant component. We consider in particular hypergraphs with higher-order triadic interactions. Higher-order triadic interactions occur when one or more nodes upregulate or downregulate a hyperedge. For instance, enzymes regulate chemical reactions involving multiple reactants. Here we propose and investigate higher-order triadic percolation on hypergraphs showing that the giant component can have a non-trivial dynamics. Specifically, we show that the fraction of nodes in the giant component undergoes a route to chaos in the universality class of the logistic map. In hierarchical higher-order triadic percolation, we extend this paradigm in order to treat hierarchically nested higher-order triadic interactions. We demonstrate the non-trivial effects of their increased combinatorial complexity on the critical phenomena and the dynamical properties of the process. Finally, we consider other generalizations of the model studying the effect of adopting interdependencies and node regulation instead of hyperedge regulation. The comprehensive theoretical framework presented here sheds light on possible scenarios for climate networks, biological networks, and brain networks, where the hypergraph connectivity changes over time.
@@ I Introduction
Higher-order networks [1, 2, 3, 4, 5, 6] capture the many-body interactions among two or more nodes present in a large variety of complex systems, ranging from the brain to biochemical reaction networks. Higher-order networks display a very rich interplay between topology and dynamics [1, 2, 7] which is leading to significant progress in network theory. Notably, higher-order interactions affect synchronization [8, 9, 10, 11, 12], percolation [13, 14, 15, 16, 17, 18, 19, 20, 21], epidemic spreading [22, 23, 24], and random walks and diffusion [25, 26].
Triadic interactions are a general type of higher-order interaction in which one or more nodes regulate the activity of a link between two other nodes. This type of interaction is widely observed in nature such as brain networks, climate networks, ecological networks, and biochemical reaction networks [27, 28, 29, 30, 31]. In the brain, the glia inhibit or facilitate the synaptic interactions between pairs of neurons. In climate networks of extreme rainfall events, triadic interactions can be used to explain the situations in which the long-range synchronized rainfall events, represented by links of the network, are modulated by large-scale patterns, such as Rossby waves [31]. In biochemical reaction networks, chemical reactions are controlled by enzymes, whose activity can be regulated by other reactions. Recently, there has been a growing interest in triadic interactions from the theoretical side. Importantly, triadic interactions have been recently shown to give rise to a paradigmatic change in the theory of percolation, according to the recently proposed triadic percolation [13, 32] model. Furthermore, it has been shown that triadic interactions affect learning [33, 34], signal processing, and network dynamics [35, 36].
@@
Percolation [38, 39, 40, 41] is one of the most fundamental critical phenomena defined on networks and has been extensively used to characterize the robustness of networks [42]. Percolation on simple networks is characterized by a second-order phase transition at which the emergence of the giant component can be observed when nodes or links are randomly damaged. In the last decades important progress in network theory has demonstrated that more general percolation problems, including k-core percolation [43], and interdependent percolation on multiplex networks [44, 45, 46, 47, 48] can display discontinuous hybrid transitions, higher-order critical points [49] and exotic phase diagrams [50]. In all these problems the size of the giant component is monitored after a (random) damage is inflicted to the nodes or to the links of the network, and after the eventual cascade of failure events triggered by the initial perturbation reaches a steady state.
@@
Triadic percolation [13, 32], however, demonstrates that percolation can become a fully-fledged dynamical process in which the giant component never reaches a steady state. If the triadic interactions are defined on top of a random network [13], the giant component becomes in general time-dependent, and its relative size, which defines the order parameter of the model, can undergo a route to chaos in the logistic map universality class [51, 52]. Moreover, triadic percolation on spatial networks leads to topological changes in the structure of the giant component generating non-trivial spatio-temporal patterns, as shown in Ref. [32].
@@
In this work, we investigate higher-order triadic percolation (HOTP) on hypergraphs. We reveal that on hypergraphs, the HOTP is a fully-fledged dynamical process whose order parameter displays period-doubling and a route to chaos when both positive and negative regulations are present with non-zero probability. Thus the phase diagram of HOTP reduces in this scenario to an orbit diagram that we hereby prove to be in the universality class of the logistic map.
@@
The critical properties of HOTP are highly non-trivial. Here we characterize them in important limiting cases in which there are exclusively positive or exclusively negative regulatory interactions, characterizing their discontinuous hybrid transitions, period-2 bifurcations, and continuous transitions. The critical conditions are theoretically predicted and numerically validated on random hypergraphs with triadic interactions. When both positive and negative regulations are present, we indicate the conditions for observing unusual orbit diagrams with a reentrant collapsed phase. In this case, we observe a phase transition between a collapsed state with a null giant component and an active phase with a non-zero and time-varying giant component while the probability of down-regulating the hyperedges increases. By numerically testing HOTP on real-world hypergraph data we observe that this dynamical behavior of HOTP remains qualitatively unchanged when HOTP occurs on non-random hypergraph architectures.
@@
Triadic interactions can also be hierarchically nested as recognized in the context of ecological networks in Ref. [29]. In order to study the effects of hierarchical regulation, we propose and study hierarchical higher-order triadic percolation (HHOTP). In this scenario, the regulatory interaction between a node and a hyperedge is controlled by other regulatory interactions and so on. We reveal that the dynamics of HHOTP is significantly different from the dynamics of HOTP due to the combinatorial complexity of their HHOTIs. Indeed as long as the hypergraphs contain negative HHOTIs with a non-vanishing probability, HHOTP displays a much richer dynamics than HOTP which can lead to period-doubling and a route to chaos also if no positive HHOTIs are present. Moreover, our theoretical derivation reveals that the route to chaos of HHOTP is no longer in the logistic map universality class.
Note that here, due to space limitation, we focus exclusively on the scenario in which nodes regulate hyperedges. However, in the future, it would be also interesting to consider HOTP in which hyperedges regulate nodes. This scenario is in line with recent results obtained in Ref. [21] where the authors consider only positive regulations.
@@
This work includes an in-depth discussion of two other interesting variations of HOTP: the interdependent higher-order triadic percolation (IHOTP), and the higher-order node dynamical percolation model (HONDP). In IHOTP, we adopt the notion of interdependent hypergraph giant component introduced in Ref. [14] which in the absence of regulation is already known to display discontinuous hybrid transition. This different nature of the underlying percolation process has dramatic effects on the dynamics of IHOTP. In particular, here we indicate the conditions under which the route-to-chaos in IHOTP is impeded. The study of HONDP allows us to discuss in detail the effect of regulating the nodes instead of the hyperedges revealing important differences with HOTP.
@@
The paper is structured as follows. In Section II, we introduce random hypergraphs with higher-order triadic interactions and with hierarchical triadic interactions; in Sections III and IV we define HOTP and HHOTP and we investigate their critical properties; in Section V, we discuss other generalizations of HOTP, including IHOTP and HONDP; finally, in Section VI we provide the concluding remarks. The paper also includes two Appendices. In the Appendix A we prove that the order parameter of HOTP undergoes a route-to-chaos in the universality class of the logistic map. In Appendix B we discuss HOTP and HONDP in the presence of partial regulation, i.e. where hyperedges (for HOTP) or nodes (for HONDP) are regulated with probability $11-\rho_{0}<$ 1.
@@ II.1 Structural hypergraphs
In this work, we consider exclusively random hypergraphs with hyperdegree distribution $P(k)$ and hyperedge cardinality distribution $Q(m)$. We denote with $\langle{k}\rangle=\sum_{k}kP(k)$ the average hyperdegree of the nodes of the hypergraph and with $\langle{m}\rangle=\sum_{m}mQ(m)$ the average cardinality of the hyperedges of the hypergraph. We furthermore assume that their factor graph representation is locally tree-like (see for a detailed discussion of this assumption Ref. [14]). This assumption is generally met by the considered random hypergraphs under very general conditions over the distributions $P(k)$ and $Q(m)$, and applies to the case in which $P(k)$ and $Q(m)$ have finite moments as for Poisson distributions. In the following we make use of the hypergraph generating functions $G_{0}(x),G_{1}(x)$ and $G_{1,m}(x)$ defined as
@@
The considered random hypergraphs reduce to networks when all the hyperedges have cardinality $m=2$, (i.e. edges). In terms of the cardinality distribution $Q(m)$, this happens when $Q(m)=\delta_{m,2}$ where here and in the following $\delta_{x,y}$ indicates the Kronecker delta. In this work, we provide general results for arbitrary hypergraphs but if not otherwise explicitly stated, we will perform the numerical analysis and the Monte Carlo simulations on hypergraphs having hyperedges with constant cardinality $m$ and Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$.
@@ II.2.1 Higher-order triadic interactions
Higher-order triadic interactions (HOTIs) occur when one or more nodes regulate a hyperedge. The HOTIs are signed, i.e. they can be associated with a positive or negative regulatory role. For instance, an enzyme can speed up or inhibit a chemical reaction. Here we consider the simple scenario in which the HOTIs occur between the same type of nodes (i.e. we do not distinguish between ”enzyme” and ”metabolite” in the previous example), generalizations in this direction will be considered in future works. In this framework, triadic interactions are encoded by a regulatory bipartite network $\mathcal{G}_{T}(V,E_{H},W)$ formed by the set of nodes $V$ and the set of factor nodes $E_{H}$ where each factor node represents a hyperedge of the structural hypergraph. The signed HOTIs are represented by the edges $W$ of the bipartite network $\mathcal{G}_{T}$. Specifically, each edge in $W$ indicates the existence of a (triadic) regulatory interaction from a node in $V$ to a hyperedge in $E_{H}$ (see Fig. 1 (a) and (b)). We call a node a positive (or negative) regulator of a hyperedge if the node regulates the hyperedge positively (or negatively). Note that the sign of regulation is a property of the regulatory interaction but not of a node.
@@ II.2.2 Hierarchical higher-order triadic interactions
Hierarchical higher-order triadic interactions (HHOTI) can be described by a multilayer regulatory network formed by $L$ layers, where each layer $\mu\in\{1,2,\ldots L\}$ is a bipartite network $\mathcal{G}_{T}^{[\mu]}$. The first layer is formed by nodes in $V$ connected to factor nodes in $E_{H}$, i.e. $\mathcal{G}_{T}^{[1]}=(V,E_{H},W^{[1]})$. Thus the first layer reduces to the fundamental HOTI defined in the previous paragraph. The second layer indicates the set of regulatory interactions $W^{[2]}$ of the regulatory interactions in $\mathcal{G}_{T}^{[1]}$, the third layer indicates the regulatory interactions $W^{3}$ of the regulatory interactions in layer 2 and so on (see Fig. 1 (e)). Thus we can describe these further layers with $\mu>1$ as bipartite networks $\mathcal{G}_{T}^{[\mu]}=(V,W^{[\mu-1]},W^{[\mu]})$ fully determining the regulatory interactions $W^{[\mu]}$ between the nodes in $V$ and the factor nodes representing the regulatory interactions $W^{[\mu-1]}$ at the previous layer.
@@
In this work, we will focus mostly on numerical results obtained for distributions $\hat{P}_{[\mu]}(\kappa^{\pm})$ independent of the layer $L$, i.e. $\hat{P}_{[\mu]}(\kappa^{\pm})=\hat{P}(\kappa^{\pm})$ with $\hat{P}(\kappa^{\pm})$ being a Poisson distribution with average $c^{\pm}$.
@@ III.1 Higher-order triadic percolation
Here we formulate higher-order triadic percolation (HOTP) that generalizes triadic percolation, proposed for networks in Ref. [13], to hypergraphs. In this framework, nodes are active if they belong to the hypergraph giant component, while hyperedges are upregulated or downregulated according to the activity of their regulator nodes and stochastic noise. The hypergraph giant component considered here is the largest extensive connected component whose nodes and hyperedges satisfy the following self-consistent and recursive conditions [14].
@@
Given a random hypergraph with hyperdegree distribution $P(k)$ and hyperedge cardinality distribution $Q(m)$ defined in Sec. II.1, at each time $t$, Step 1 implements hyperedge percolation [14] where the probability that a hyperedge is intact is given by $p_{H}^{(t-1)}$. This process determines the fraction of nodes in the hypergraph giant component $R^{(t)}$ at time $t$, given by
@@
where $\hat{S}^{(t)}$ is the probability that starting from a random node and choosing one of its hyperedges at random we reach a hyperedge in the hypergraph giant component. The probability $\hat{S}^{(t)}$ together with the probability $S^{(t)}$ that starting from a random hyperedge and choosing one of its nodes at random we reach a node in the hypergraph giant component, obey the following self-consistent set of equations
@@
Thus Eqs. (6) and (7) together with Eq. (5) determine $R^{(t)}$ starting from the knowledge of $p_{H}^{(t)}$ and implements Step 1. Note that this step differs from Step 1 of standard triadic percolation [13], as here we define a node active only if it belongs to the hypergraph giant component while in the standard triadic percolation we define a node active only if it belongs to the standard giant component of the network under consideration.
Step 2 requires the formulation of an additional equation determining the probability $p_{H}^{(t)}$ that the hyperedges are intact given the probability $R^{(t)}$ that the nodes are active, (i.e. they belong to the hypergraph giant component at Step 1). On the hypergraphs with random interactions defined in Sec. II.2.1, $p_{H}^{(t)}$ obeys the following dynamical equations:
@@
where $G_{0}^{\pm}(x)$ are the generating functions defined in Eq. (3). Note that Eq. (8) is the same equation used in Ref. [13] to define Step 2 of standard triadic percolation, however here $p_{H}^{(t)}$ indicates the probability that a hyperedge is intact while in standard triadic percolation it indicates the probability that an edge is intact.
From this definition of higher-order triadic percolation, it is clear that this dynamical process reduces to standard triadic percolation [13] if $P(m)=\delta_{m,2}$, i.e. if the hypergraph reduces to a network.
Figure 2: Time series (panels (a)-(d)) of the order parameter $R$ of HOTP and their corresponding Cobwebs (panels (e)-(h)) are shown to display a variety of dynamical behavior: a steady state (panels (a) and (e)), period-2 oscillations (panels (b) and (f)), period-4 oscillations (panel (c) and (g)) and chaotic dynamics (panel (d) and (h)). The time series are obtained by Monte Carlo simulations on hypergraphs of $N=10^{6}$ nodes with HOTIs. The structural hypergraph has hyperedges of fixed cardinality $m=10$ and Poisson hyperdegree distribution $P(k)$ with an average $c=10$. The degree distributions $\hat{P}(\kappa^{\pm})$ are Poisson with an average regulatory positive and negative degree $c^{+}=3$ and $c^{-}=5.5$ respectively. In the Cobweb, the blue curves denote function $R=f_{m}(p_{H})$ and the orange curves denote function $p_{H}=g_{p}(R)$ as defined in Eq. (9) and Eq. (10). The time series and the Cobweb plots are obtained for $p=0.27$ (panels (a) and (e)), $p=0.35$ (panels (b) and (f)), $p=0.68$ (panels (c) and (g)) and $p=0.88$ (panels (d) and (h)).
Figure 3: Examples of the map $h_{m,p}(R)$ encoding for the dynamics of HOTP when the HOTIs are exclusively positive (panel (a)); when the HOTIs are exclusively negative (panel (b)) and when the hypergraphs include both positive and negative HOTIs with non-vanishing probability (panel (c)). The map is unimodal only the scenario (c), thus we can observe a route to chaos of the order parameter only in the presence of both positive and negative HOTIs. For all three panels, the structural networks have constant hyperedge cardinality $m=3$ and Poisson hyperdegree distribution with an average $c=4$. Moreover the data ar obtained for $\hat{P}^{\pm}(\kappa)$ with average degree $c^{\pm}$ given by $c^{+}=2,c^{-}=0$ (panel (a)); $c^{+}=0,c^{-}=2$ (panel (b)).; $c^{+}=2,$ $c^{-}=1.5$ (panel (c)).